The demand and supply equations for a commodity (in a free market) are given as;





Qd = 10 - 2p





As = 4p - 8





a. Given that p is in naira, Qd and Qd are in kg. Determine





(a) the equilibrium price and





(b) the equilibrium quantity





b. (i) if the price (p) were to be N4.00, what will be excess supply?





(ii) if the price were fixed at N1, what will be the excess demand?

1
Expert's answer
2021-11-24T12:29:40-0500

(a)

Equilibrium price:

At equilibrium, the quantity demanded is equal to the quantity supplied.

"Q_d=Q_s"

"10-2P=4P-8"

"P=8"

Equilibrium price=8 naira.


(b)

Equilibrium quantity:

"=4(8)-8=24" Kg.